# Enhancing Pre-Trained Decision Transformers with Prompt-Tuning Bandits

### v2

@@ Abstract
Harnessing large offline datasets is vital for training foundation models that can generalize across diverse tasks. Offline Reinforcement Learning (RL) offers a powerful framework for these scenarios, enabling the derivation of optimal policies even from suboptimal data. The Prompting Decision Transformer (PDT) is an offline RL multi-task model that distinguishes tasks through stochastic trajectory prompts, which are task-specific tokens maintained in context during rollouts. However, PDT samples these tokens uniformly at random from per-task demonstration datasets, failing to account for differences in token informativeness and potentially leading to performance degradation. To address this limitation, we introduce a scalable bandit-based prompt-tuning method that dynamically learns to construct high-performance trajectory prompts. Our approach significantly enhances downstream task performance without modifying the pre-trained Transformer backbone. Empirical results on benchmark tasks and a newly designed multi-task environment demonstrate the effectiveness of our method, creating a seamless bridge between general multi-task offline pre-training and task-specific online adaptation.
Machine Learning, ICML
@@ 1 Introduction
The ability to exploit large amounts of offline data is crucial for training foundation models capable of generalizing across diverse tasks (Radford, 2018; Reed et al., ; Brohan et al., 2022). In Reinforcement Learning (RL), a seminal contribution is the Decision Transformer (DT) (Chen et al., 2021), which reframes offline RL (Levine et al., 2020) as a sequence modeling problem, thereby leveraging powerful Transformer architectures. The Prompting Decision Transformer (PDT) (Xu et al., 2022) further extends DT from single-task to multi-task settings, enabling large-scale models and generalized pre-training in offline meta- and few-shot RL (Xu et al., 2022; Mitchell et al., 2021). Analogous to prompting in Large Language Models (LLMs), PDT differentiates tasks through a stochastic trajectory prompt prepended to the context, allowing it to identify and model optimal action marginals for each task in the offline dataset.
However, PDT samples segments of prompts uniformly at random from expert demonstrations, overlooking a key limitation: not all prompts are equally informative for differentiating tasks. Surprisingly, this issue persists even in offline datasets from fully observable Markov Decision Processes (MDPs). We hypothesize that the sampling of non-informative prompts from the expert demonstrations can diminish PDT’s ability to differentiate between tasks, thereby leading to performance degradation.
To address this shortcoming, we introduce a scalable, robust, and computationally efficient bandit-based prompt-tuning method for PDT. By optimizing prompt selection, our approach boosts downstream task performance without costly weight updates to the underlying Transformer backbone. Our experiments reveal clear performance gains with the proposed method, effectively bridging the gap between pre-training and adaptation.
@@ 2.1 Problem definition: Offline multi-task RL
$\mathbf{x}=\big{(}\rho\big{)}\odot\big{(}\overbrace{\hat{r}_{t-K},\mathbf{s}_{t-K},\mathbf{a}_{t-K},\hat{r}_{t-K+1},\mathbf{s}_{t-K+1},\mathbf{a}_{t-K+1}\dots,\hat{r}_{t},\mathbf{s}_{t},\mathbf{a}_{t}}^{\omega_{K:t}:K\text{most recent transitions}}\big{)}$ (2)
@@
The goal is to learn a generalized policy, $\pi(\mathbf{s},\psi)\to\mathbf{a}$, capable of solving all tasks in $\mathcal{T}^{\text{train}}$. Here, $\psi$ serves as an auxiliary input to the policy that sufficiently describes the task $\mathcal{T}_{i}$. This input can take forms such as a task index, one-hot embedding, or task parameter, ensuring that the policy is aware of the current target task. In the case of PDT, $\psi$ is the stochastic trajectory prompt sampled from the demonstration set for the target task. For each task, the optimal generalized policy is expected to maximize the corresponding expected discounted reward objective specific to task $i$.
$J(\pi,\psi)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma_{i,t}r_{i}(\mathbf{s}_{t},\mathbf{a}_{t})\right]$ (3)
@@ 2.2 Prompting Decision Transformer
Chen et al. (2021) reframed RL as a sequence modeling problem, leveraging the capabilities of Transformer architectures to model trajectories. A trajectory in DT is represented as a sequence of triplets $(\hat{r}_{t},\mathbf{s}_{t},\mathbf{a}_{t})$, where $\hat{r}_{t}=\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}$ is the return-to-go, $\mathbf{s}_{t}$ is the state, and $\mathbf{a}_{t}$ is the action at time step $t$. DT captures the temporal and causal dependencies between states, actions, and rewards, enabling the model to predict optimal actions directly, without relying on explicit value functions or policies. This design is particularly effective in offline RL settings, where pre-collected datasets of trajectories are available for training.
@@
For a particular training task $\mathcal{T}_{i}\in\mathcal{T}^{\text{train}}$, PDT learns to model the sequence in Equation 2 by autoregressively predicting the action tokens, where $\odot$ denotes concatenation. While PDT randomly samples the $J$ segments that constitute the prompt $\rho$ from a set of expert demonstrations $\mathcal{P}_{i}$ for each task, we propose a method to optimize segments selection, aiming to enhance the downstream task performance.
@@ 2.3 Multi-Armed Bandits
Multi-Armed Bandits (MABs) provide a framework for optimizing stochastic reward functions, making them an effective tool for tasks like prompt optimization. In the standard MAB setting, at each time step $k$, the agent selects an action (or “arm”) $a_{k}\in\mathcal{A}$, where $\mathcal{A}$ is the set of available arms. Here, we use $k$ to denote bandit time steps, to avoid confusion with the MDPs time $t$. The agent then receives a stochastic reward $r_{k}\sim R(a_{k})$, with the goal of maximizing the cumulative reward $\sum_{k=1}^{K}r_{k}$ over a time horizon $K$ (Auer et al., 2002). This requires balancing the exploration of arms to gather information about their reward distributions $R(a)$ and exploitation of arms with known high rewards while minimizing cumulative regret, defined as:
$\text{Regret}(K)=\sum_{k=1}^{K}\left[\max_{a\in\mathcal{A}}\mathbb{E}[R(a)]-\mathbb{E}[r_{k}]\right]$ (4)
@@ 3 Related Work
Recently, there has been a surge in methods across various domains aimed at enhancing the performance and generalization of Transformer-based approaches through automatic prompt tuning. The main challenge lies in the computationally expensive optimization of discrete prompts, guided by stochastic feedback from evaluating a black-box target model.
@@
Prompting in LLMs. For LLMs, Chen et al. (2024) introduced InstructZero, which uses Bayesian optimization to explore low-dimensional soft prompt vectors. These vectors are then passed to an open-source LLM to generate instructions for the black-box LLM. Building on this, Lin et al. (2023) proposed INSTINCT, which replaces the Gaussian process in Bayesian optimization with a neural network surrogate, leveraging a neural bandit algorithm to enhance expressivity. Additionally, Shi et al. (2024a) demonstrated that the fixed-budget MAB framework enables learning the optimal prompt within a limited number of LLM evaluations. These methods rely on optimization in continuous spaces, followed by prompt reconstruction, whereas our approach directly operates within the DT prompt space.
Another line of research on prompt augmentation, or demonstration learning, investigates how selecting examples from a demonstration pool affects downstream performance. Liu et al. (2021) utilized sentence embeddings to choose examples aligned with the input, while Mishra et al. (2021) highlighted the benefits of combining both positive and negative examples for improved guidance. Approaches by Kumar & Talukdar (2021) and Lu et al. (2021) focused on optimizing the order of demonstrations using separator tokens and entropy-based scoring. In contrast, our method moves beyond reliance on the selection and ordering of the trajectory prompts and focuses on constructing optimal prompts by combining different segments from these trajectory prompts.
Multi-task DT. For RL, Hu et al. (2023) proposed to generate DT prompt candidates by perturbing the initial trajectory with Gaussian noise and using online or offline training to compute feedback for a ranking optimization algorithm. Similarly, the Prompt Diffuser (Hu et al., 2024) frames instruction optimization as a conditional generative modeling task, generating prompts starting from random noise. In contrast, our method focuses on constructing prompts from expert demonstrations, as noise-based approaches are not suitable in certain settings, such as discrete spaces. Wang et al. (2024) proposed hierarchical prompting that provides context-specific guidance through two levels of prompting: one that encapsulates task-specific information and another that uses a set of demonstration segments to guide the rollouts. Hyper-decision transformer (Xu et al., 2023) augmented the base DT with a hyper-network module to achieve adaptation to unseen novel tasks. However, both of those methods rely mainly on the presence of larger amounts of demonstration data, while our approach requires only a few demonstrations from which we learn to select the best prompt.
@@
LLM for RL. At the intersection of LLM and RL domains, Yang & Xu (2024) and the closely related LaMo method (Shi et al., 2024b) proposed leveraging a pre-trained LLM as an initializer for PDT, harnessing rich linguistic knowledge to boost performance on unseen tasks. In another work, Zheng et al. (2024) introduced an approach to decompose the prompt into cross-task and task-specific components, ensuring more robust test-time adaptation to unseen tasks. Furthermore, the model is initialized by incorporating parameters from a pre-trained language model to provide it with prior knowledge. Compared to those, our approach avoids reliance on an additional large model, sidestepping the fine-tuning and scalability challenges inherent in such methods.
Figure 1: Overview of our inference time bandit-based prompt tuning for multi-task learning with prompting decision transformer (PDT). Each $z_{i}$ represents a triplet $(\hat{r}_{i},\mathbf{s}_{i},\mathbf{a}_{i})$. The bandit explores the demonstration dataset $\mathcal{P}_{i}$ for the current task $i$ to find the segments $\tilde{\tau}$ that induce the best prompt. The online return $G_{k}$ achieved by the underlying PDT model at round $k$ and using prompt $\rho_{k}$ serves as reward for the bandit. For simplicity in our illustration, we set $H=2$.
@@ 4 Method
In this section, we present an inference time prompt-tuning method for offline multi-task RL.
Given a dataset $\mathcal{D}=\{\mathcal{D}_{0},\dots,\mathcal{D}_{n}\}$ of trajectories for $n$ training tasks $\mathcal{T}^{\text{train}}$, we utilize the original PDT method to learn the generalized policy described in Equation 3. Details of the algorithm and training process are provided by Xu et al. (2022).
An optimal PDT model $\pi^{*}(\mathbf{x};\theta)$ is assumed to be trained until convergence on $\mathcal{D}$. After training, the goal is to evaluate and enhance the model’s performance on a specific task $\mathcal{T}_{i}$, either from the set of holdout test tasks $\mathcal{T}^{\text{test}}$ or from the training set $\mathcal{T}^{\text{train}}$. We assume access to a small set of demonstrations $\mathcal{P}_{i}$ (not necessarily of expert quality) for the target task, which serve as a source for sampling prompts, as well as a simulator of the corresponding $\mathcal{M}_{i}$ to perform online evaluations. Our primary objective is to identify the stochastic trajectory prompt in $\mathcal{P}_{i}$ that maximizes performance when applied to the pre-trained generalized policy.
@@ 4.1 Bandit-based Prompt Tuning
We propose leveraging a bandit-based approach to efficiently identify the optimal trajectory prompt. A naïve implementation of bandit-based prompt tuning would involve deploying a bandit with one arm for each possible prompt $\rho$ constructible from $\mathcal{P}_{i}=(\tau_{1},\dots,\tau_{n})$, i.e. considering all possible combinations of segments among demonstration trajectories. This approach, however, scales extremely poorly, as the number of prompts (and consequently, the size of the bandit problem) grows linearly with the total number of transitions in $\mathcal{P}_{i}$ and combinatorially with $J$, the number of segments in each prompt. Additionally, treating each prompt as a separate and independent arm disregards prompt similarity, leading to inefficient sample complexity.
We propose optimizing the selection of segments $\tilde{\tau}$ that form the stochastic trajectory prompt $\rho=(\tilde{\tau}_{1},\dots,\tilde{\tau}_{J})$ using a contextual MAB algorithm. Unlike traditional contextual MAB settings, where contexts $\mathbf{c}_{k}$ are revealed incrementally at each bandit time step $k$, our framework assumes that the entire dataset $\mathcal{P}_{i}$ is available upfront.
As illustrated in Figure 1, at each bandit step $k$, the bandit selects a prompt $\rho_{k}$, balancing exploration and exploitation within the prompt space. In each round, the bandit selects $J$ segments to construct the trajectory prompt $\rho_{k}$ which is then passed to the pre-trained PDT $\pi^{*}(\mathbf{x}_{k};\theta)$. The PDT is rolled out in $\mathcal{M}_{i}$ using the selected prompt, and the resulting performance is logged as $G_{i}^{k}=\sum_{t=0}^{T}r_{i}(\mathbf{s}_{t},\mathbf{a}_{t})\mid\mathbf{a}_{t}\sim\pi^{*}(\mathbf{x}_{k};\theta)$. The input to the PDT at each MDP step $t$ is given by $\mathbf{x}_{k}=\rho_{k}\odot\omega_{K:t}$ (see Eq. 2), where the trajectory prompt $\rho_{k}$ remains fixed throughout the rollout, and the MDP transitions $\omega_{K:t}$ are updated dynamically after each step. The process of constructing the prompt $\rho_{k}$ through the bandit mechanism is discussed in detail in the next section.
The performance of rollout $k$, $G_{i}^{k}$, serves as the reward from the bandit’s perspective, and the tuple $\langle\rho_{k},G_{i}^{k}\rangle$ is stored for training the reward model. The pseudocode of the proposed method is provided in Algorithm 1.

### v3

# Prompt Tuning Decision Transformers with Structured and Scalable Bandits

@@ Abstract
Prompt tuning has emerged as a key technique for adapting large pre-trained Decision Transformers (DTs) in offline Reinforcement Learning (RL), particularly in multi-task and few-shot settings. The Prompting Decision Transformer (PDT) enables task generalization via trajectory prompts sampled uniformly from expert demonstrations – without accounting for prompt informativeness. In this work, we propose a bandit-based prompt-tuning method that learns to construct optimal trajectory prompts from demonstration data at inference time. We devise a structured bandit architecture operating in the trajectory prompt space, achieving linear rather than combinatorial scaling with prompt size. Additionally, we show that the pre-trained PDT itself can serve as a powerful feature extractor for the bandit, enabling efficient reward modeling across various environments. We theoretically establish regret bounds and demonstrate empirically that our method consistently enhances performance across a wide range of tasks, high-dimensional environments, and out-of-distribution scenarios, outperforming existing baselines in prompt tuning.
@@ 1 Introduction
The ability to exploit large amounts of offline data is crucial for training foundation models capable of generalizing across diverse tasks (Radford, 2018; Reed et al., 2022; Brohan et al., 2022). In Reinforcement Learning (RL), a seminal contribution is the Decision Transformer (DT) (Chen et al., 2021), which reframes offline RL (Levine et al., 2020) as a sequence modeling problem, thereby unlocking powerful Transformer architectures for offline RL. DT is particularly well-suited for offline RL because it sidesteps the well-known instabilities of temporal-difference learning with function approximation (Sutton & Barto, 2018), which are exacerbated under distribution shift in the offline setting, by replacing discounted, moving-target value estimates with stationary return-to-go conditioning, autoregressive sequence modeling, and causal attention mechanisms (Chen et al., 2021). The Prompting Decision Transformer (PDT) (Xu et al., 2022) further extends DT from single-task to multi-task settings, enabling large-scale models and generalized pre-training in offline multi-task and few-shot RL (Xu et al., 2022; Mitchell et al., 2021). Analogous to prompting in Large Language Models (LLMs), PDT differentiates tasks through a stochastic trajectory prompt prepended to the context, allowing it to identify and model optimal action marginals for each task in the offline dataset. This makes PDT particularly appealing for offline multi-task RL, as it avoids additional learning on the downstream task while enabling efficient and robust adaption of the model’s behavior at inference time, solely through adjustments of the prompt.
However, PDT samples prompts uniformly, overlooking that not all prompts are equally informative, even in fully observable MDP. We hypothesize that the sampling of non-informative prompts from expert demonstrations can diminish PDT’s ability to differentiate between tasks, thereby leading to performance degradation.
While improving on PDT’s uninformed prompt-sampling strategy, prior works on PDT prompt-tuning suffer from key limitations. The approach by Yuan et al. (2024) replaces trajectory prompts with less-expressive goal-conditioning and relies on hindsight relabeling. The generative approaches by Hu et al. (2023, 2024) are not applicable in discrete settings and don’t adhere to causal relationships between prompt tokens. Furthermore, all of these works treat prompts as flat, unstructured inputs and operate directly on MDP modalities, which leads to poor scaling with prompt size and state- and action-space sizes.
To address these shortcomings, we introduce a scalable, robust, and computationally efficient bandit-based prompt-tuning method for PDT. By exploiting the structure of the prompt space, our method reduces the complexity of prompt selection from combinatorial to linear in the prompt size. Moreover, we show how to leverage the pre-trained PDT as a feature extractor to obtain compact prompt representations that enable efficient deployment even in high-dimensional (e.g. pixle-based) settings. By optimizing prompt selection, our approach boosts downstream task performance without costly weight updates to the underlying Transformer backbone. Our experiments reveal clear performance gains with the proposed method, effectively bridging the gap between pre-training and adaptation.
@@ 2.1 Problem definition: Offline multi-task RL
The goal is to learn a generalized policy, $\pi(\mathbf{s},\psi)\to\mathbf{a}$, capable of solving all tasks in $\mathcal{T}^{\text{train}}$. Here, $\psi$ serves as an auxiliary input to the policy that sufficiently describes a task from $\mathcal{T}$. In the case of PDT, $\psi$ is the stochastic trajectory prompt sampled from the demonstration set for the target task. For each task, the optimal generalized policy is expected to maximize the corresponding expected discounted reward objective specific to task $i$.
$J(\pi,i)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma_{i}^{t}r_{i}(\mathbf{s}_{t},\mathbf{a}_{t})\right]$ (1)
@@ 2.2 Prompting Decision Transformer
Chen et al. (2021) reframed RL as a sequence modeling problem, leveraging the capabilities of Transformer architectures to model trajectories. A trajectory in DT is represented as a sequence of triplets $(\hat{r}_{t},\mathbf{s}_{t},\mathbf{a}_{t})$, where $\hat{r}_{t}=\sum_{t^{\prime}=t}^{T}r_{t^{\prime}}$ is the return-to-go, $\mathbf{s}_{t}$ is the state, and $\mathbf{a}_{t}$ is the action at time step $t$. DT captures the temporal and causal dependencies between states, actions, and rewards, enabling the model to predict optimal actions directly, without relying on explicit value functions or policies.
@@
$\mathbf{x}$ $=\big(\rho\big)\odot\big(\overbrace{\hat{r}_{t-L},\mathbf{s}_{t-L},\mathbf{a}_{t-L},\hat{r}_{t-L+1},\mathbf{s}_{t-L+1},\mathbf{a}_{t-L+1}\dots,\hat{r}_{t},\mathbf{s}_{t},\mathbf{a}_{t}}^{\omega_{L:t}:L\text{ most recent transitions}}\big)$ (3)
For a particular training task $\mathcal{T}_{i}\in\mathcal{T}^{\text{train}}$, PDT learns to model the sequence in Equation 3 by autoregressively predicting the action tokens, where $\odot$ denotes concatenation. While PDT randomly samples the $J$ segments that constitute the prompt $\rho$ from a set of expert demonstrations $\mathcal{P}$, we propose to optimize segment composition to enhance the downstream task performance with a Multi-Armed Bandit approach.
@@ 2.3 Multi-Armed Bandits
Multi-Armed Bandits (MABs) provide a framework for optimizing stochastic reward functions, making them an effective tool for tasks like prompt tuning. In the standard MAB setting, at each time step $k$, the agent selects an action (or “arm”) $a_{k}\in\mathcal{A}$, where $\mathcal{A}$ is the set of available arms. Here, we use $k$ to denote bandit time steps, to avoid confusion with the MDPs time $t$. The agent then receives a stochastic reward $r_{k}\sim R(a_{k})$, with the goal of maximizing the cumulative reward $\sum_{k=1}^{K}r_{k}$ over a time horizon $K$ (Auer et al., 2002). This requires balancing the exploration of arms to gather information about their reward distributions $R(a)$ and exploitation of arms with known high rewards while minimizing cumulative regret, defined as:
$\mathrm{Regret}(K)=\sum_{k=1}^{K}\left[\max_{a\in\mathcal{A}}\mathbb{E}[R(a)]-\mathbb{E}[r_{k}]\right]$ (4)
@@ 3 Related Work
Recently, there has been a surge in methods across various domains aimed at enhancing the performance and generalization of Transformer-based approaches through automatic prompt tuning.
@@
Prompting in LLMs. For LLMs, Chen et al. (2024) introduced InstructZero, which uses Bayesian optimization to explore low-dimensional soft prompt vectors. These vectors are then passed to an open-source LLM to generate instructions for the black-box LLM. Building on this, Lin et al. (2023) proposed INSTINCT, which replaces the Gaussian process in Bayesian optimization with a neural network surrogate, leveraging a neural bandit algorithm to enhance expressivity. Additionally, Shi et al. (2024a) demonstrated that the fixed-budget MAB framework enables learning the optimal prompt within a limited number of LLM evaluations. These methods rely on optimization in continuous spaces, followed by prompt reconstruction, whereas our approach directly operates within the PDT prompt space.
Multi-task DT. For RL, Hu et al. (2023) proposed generating prompt candidates for PDT by perturbing initial trajectories with Gaussian noise and applying online or offline feedback to guide a ranking-based optimization. Prompt Diffuser (Hu et al., 2024) similarly frames instruction optimization as a conditional generative modeling task, synthesizing prompts from random noise. In contrast, our method constructs prompts directly from expert demonstrations, which is particularly beneficial in settings like discrete state and action spaces, where noise-based approaches are less effective.
Wang et al. (2024a) introduced hierarchical prompting, using two levels of context: one for encoding task-specific information and another for guiding rollouts via demonstration segments. Hyper-Decision Transformer (Xu et al., 2023) augments DT with a hyper-network for adaptation to novel tasks, but both methods rely on large quantities of demonstration data. Our method, by contrast, selects high-quality prompts from only a few expert trajectories. Other approaches focus on representation learning or adaptation, e.g., Wang et al. (2024b) focus on disentangled world models, while Xie et al. (2023) propose latent-variable conditioned RL and fine-tuning. Our approach avoids updating the transformer backbone and relies solely on test-time prompt optimization.
Finally, MGPO (Yuan et al., 2024) proposes a general framework for online prompt tuning. While sharing a similar goal, we provide a method to identify the best prompt from a demonstration dataset, while MGPO iteratively refines the prompt with data collected during online rollouts.
@@
LLM for RL. At the intersection of LLM and RL domains, Yang & Xu (2024) and the closely related LaMo method (Shi et al., 2024b) proposed leveraging a pre-trained LLM as an initializer for PDT, harnessing rich linguistic knowledge to boost performance on unseen tasks. In another work, Zheng et al. (2024) introduced an approach to decompose the prompt into cross-task and task-specific components, ensuring more robust test-time adaptation to unseen tasks. Furthermore, the model is initialized by incorporating parameters from a pre-trained language model to provide it with prior knowledge. Compared to those, our approach avoids reliance on a dedicated LLM, sidestepping the fine-tuning and scalability challenges inherent in such methods.
@@ 4 Method
We now present our inference time prompt-tuning bandit method for offline multi-task RL. Given a dataset $\mathcal{D}=\{\mathcal{D}_{0},\dots,\mathcal{D}_{n}\}$ of trajectories for $n$ training tasks $\mathcal{T}^{\text{train}}$, we utilize the original PDT method to learn the generalized policy described in Equation 1. Details of the PDT algorithm and training process are provided by Xu et al. (2022).
An optimal PDT model $\pi^{*}(\mathbf{x};\theta)$ is assumed to be trained until convergence on $\mathcal{D}$. After training, the goal is to evaluate and enhance the model’s performance on a specific task $\mathcal{T}_{i}$, either from the set of holdout test tasks $\mathcal{T}^{\text{test}}$ or from the training set $\mathcal{T}^{\text{train}}$. We assume access to a small set of demonstrations $\mathcal{P}_{i}$ for the target task, which serve as a source for sampling prompts, as well as a simulator of the corresponding $\mathcal{M}_{i}$ to perform online evaluations. Our primary objective is to identify the stochastic trajectory prompt constructible from $\mathcal{P}_{i}$ that maximizes performance when applied to the pre-trained generalized policy.
@@ 4.1 Bandit-based Prompt Tuning
Figure 1: Overview of our bandit-based prompt-tuning method for multi-task learning with PDT. Each $z_{i}$ represents a triplet $(\hat{r}_{i},\mathbf{s}_{i},\mathbf{a}_{i})$, each $\tilde{\tau}$ represents a prompt segment, each $\tau$ represents a demonstration trajectory. The bandit explores the demonstration dataset $\mathcal{P}_{i}$ for the current task $i$ to find the best prompt $\rho^{*}=(\tilde{\tau}_{1}^{*},\dots,\tilde{\tau}_{J}^{*})$. The online return $G_{k}$ achieved by the underlying PDT model at round $k$ and using prompt $\rho_{k}$ serves as a reward for the bandit.
We propose leveraging a bandit-based approach to efficiently identify the optimal trajectory prompt. A naïve implementation of bandit-based prompt tuning would involve deploying a bandit with one arm for each possible prompt $\rho$ constructible from $\mathcal{P}_{i}$, i.e. considering all possible combinations of segments among demonstration trajectories. This approach, however, scales extremely poorly, as the number of prompts (and consequently, the size of the bandit problem) grows linearly with the total number of transitions in $\mathcal{P}_{i}$ and combinatorially with $J$, the number of segments in each prompt. Additionally, treating each prompt as a separate and independent arm disregards prompt similarity, leading to inefficient sample complexity. Thus, we propose to optimize the selection of segments $\tilde{\tau}$ that form the stochastic trajectory prompt $\rho=(\tilde{\tau}_{1},\dots,\tilde{\tau}_{J})$ with a special contextual MAB algorithm.
As illustrated in Figure 1, at each round $k$, the bandit selects a prompt $\rho_{k}$ consisting of $J$ segments, balancing exploration and exploitation in the prompt space. The process of constructing the prompt $\rho_{k}$ through the bandit mechanism is discussed in detail in the next section. We then perform the $k$-th rollout of the PDT $\pi^{*}(\mathbf{x};\theta)$ in $\mathcal{M}_{i}$, using the selected prompt $\rho_{k}$ and the most recent transitions, as per Equation 3. The resulting performance $G_{i}^{k}=\sum_{t=0}^{T}r_{i}(\mathbf{s}_{t},\mathbf{a}_{t})\mid\mathbf{a}_{t}\sim\pi^{*}(\mathbf{x}_{k};\theta)$ serves as reward from the bandit’s perspective and is stored together with the prompt $\rho_{k}$ for training the bandit. The algorithmic pseudocode for prompt-tuning with our bandit is provided in the Algorithm 1 in Appendix A.